Interest question, please help
$5400 is invested, part of it at 10% and part of it at 7%. For a certain year the total yield is $468. How much was invested at each rate? Thanks for your help.

Answer by stanbon(57214)

Interest question, please help
$5400 is invested, part of it at 10% and part of it at 7%. For a certain year the total yield is $468. How much was invested at each rate?
-----------------
Quantity Equation: x + y = 5400 dollars
Interest Equation:0.10x + 0.07y = 469 dollars
-------------------------------------------
Multiply thru 1st by 10
Multiply thru 2nd by 100
--------------------------------
10x + 10y = 54000
10x + 7y = 46900
---------------------------
Subtract and solve for "y":
3y = 7100
y = $2366.67 (amt. invested at 7%)
---
Solve for "x":
x+ y = 5400
x = 5400-2366.67
x = $2633.33 (amt. invested at 10%)
=======================================
Cheers,
Stan H.
=============